# Manifolds of mappings associated with real-valued functions spaces

### v1

@@ 1 Introduction
For fixed $m,n\in\mathbb{N}$, we consider a $m$-dimensional compact smooth manifold with corners $M$ and $N$ be a $n$-dimensional smooth manifold without boundary. We consider a basis of the topology $\mathcal{U}$ of the set $[0,\infty)^{m}$ satisfying suitable properties (see Definition 2.1). Suppose that for each open set $U\in\mathcal{U}$, an integral complete locally convex space $\mathcal{F}(U,\mathbb{R})$ of bounded, continuous real-valued functions are given. Then for each finite-dimensional real vector space $E$, the set of maps $\mathcal{F}(U,E)$ can be defined in a natural way. If certain axioms are satisfied (see Definition 2.5), we say that the family $(\mathcal{F}(U,E))_{U\in\mathcal{U}}$ is suitable for global analysis. As direct consequence of the case where $M$ is a smooth manifold without boundary (see [15]), it can be shown that one can define a locally convex space $\mathcal{F}(M,E)$. Moreover, we can also define the set of mappings $\mathcal{F}(M,N)$ of $N$-valued functions on the manifold with corners $M$.
@@
Let $N$ be a smooth manifold and $\pi_{TN}:TN\to N$ its tangent bundle. A local addition is a smooth map ${\Sigma:\Omega\to N}$ defined on a open neightborhood $\Omega\subseteq TN$ of the zero-section ${0_{N}:=\{0_{p}\in T_{p}N:p\in N\}}$ such that $\Sigma(0_{p})=p$ for all $p\in N$ and the image $\Omega^{\prime}:=\big{(}\pi_{TN},\Sigma\big{)}(\Omega)$ is open in $N\times N$ and the map $\theta_{N}:=(\pi_{TN},\Sigma):\Omega\to\Omega^{\prime}$ is a $C^{\infty}$-diffeomorphism.
@@
$\Gamma_{\mathcal{F}}(\eta):=\{\sigma\in\mathcal{F}(M,TN):\pi_{TN}\circ\sigma=\gamma\}$
@@
Let $M$ be a $m$-dimensional compact smooth manifold with corners, $N_{1}$ and $N_{2}$ be $n$-dimensional smooth manifold which admits local addition $(\Omega_{1},\Sigma_{1})$ and $(\Omega_{2},\Sigma_{2})$ respectively. If $f:N_{1}\to N_{2}$ is a $smooth$ map, then the map
@@
Let $M$ be a $m$-dimensional compact smooth manifold with corners, $N$ be a $n$-dimensional smooth manifold which admits a local addition and $\pi_{TN}:TN\to N$ its tangent bundle. Then the map
$\mathcal{F}(M,\pi_{TN}):\mathcal{F}(M,TN)\to\mathcal{F}(M,N),\quad\tau\mapsto\pi_{TN}\circ\tau$
@@ 2 Preliminaries
Let $m\in\mathbb{N}$ fixed, a set $\mathcal{U}$ of open subsets of product set $[0,\infty)^{m}$ will be called a good collection of open subsets if the following condition are satisfied:
@@
    $\mathcal{U}$ is a basis for the topology of $[0,\infty)^{m}$.
@@
    If $U\in\mathcal{U}$ and $K\subseteq U$ is a compact non-empty subset, then there exists $V\in\mathcal{U}$ with compact closure $\overline{V}$ in $[0,\infty)^{m}$ such that $K\subseteq V$ and $\overline{V}\subseteq U$.
@@
    If $\phi:U\to V$ is a $C^{\infty}$-diffeomorphism between open subsets $U$ and $V$ of $[0,\infty)^{m}$ and $W\in\mathcal{U}$ is a relatively compact subset of $U$, then $\phi(W)\in\mathcal{U}$.
@@
Let $U$ be a open subset of $[0,\infty)^{m}$, we write $BC(U,\mathbb{R})$ for the vector space of all bounded continuous functions $f:U\to\mathbb{R}$ endowed with the supremum norm $\lVert\cdot\lVert_{\infty}$.
@@
Let $\mathcal{U}$ be a good collection of open subsets of $[0,\infty)^{m}$. For $U\in\mathcal{U}$, the vector subspace $\mathcal{F}(U,\mathbb{R})$ of $BC(U,\mathbb{R})$ will denote a integral complete locally convex space such that the inclusion map $\mathcal{F}(U,\mathbb{R})\to BC(U,\mathbb{R})$ is continuous.
@@
    $f_{*}:\mathcal{F}(U,E)\to F(V,F),\quad\gamma\mapsto f\circ(\text{id}_{V},\gamma|_{V})$
@@
    Pullback Axiom : Let $U$ be an open subset of $[0,\infty)^{m}$ and $V,W\in\mathcal{U}$ such that $W$ has compact closure contained in $U$. Let $\Theta:U\to V$ be a smooth diffeomorphism. Then $\gamma\circ\Theta|_{W}\in\mathcal{F}(W,E)$ for all $\gamma\in\mathcal{F}(V,E)$ and
@@
Following [15, Remark 3.5], if $\mathcal{U}$ is a good collection of open subsets of $[0,\infty)^{m}$ and $\left(\mathcal{F}(U,\mathbb{R})\right)_{U\in\mathcal{U}}$ is a family of locally convex space suitable for global analysis, then we have the following results.
@@
which extends functions by 00 is a linear isometry. ∎
@@
[…]
@@
Let $U$ be a open subset of $[0,\infty)^{m}$ and $E$ be a finite-dimensional real vector space. We let $\mathcal{F}_{\text{loc}}(U,E)$ be the set of all function $\gamma:U\to E$ such that for each $V\in\mathcal{U}$ which is relatively compact in $U$ we have $\gamma|_{V}\in\mathcal{F}(V,E)$.
@@
Let $E$ be a finite-dimensional vector space. If $U$ and $V$ are open subsets of $[0,\infty)^{m}$ such that $V\subseteq U$, then $\gamma|_{V}\in\mathcal{F}_{\text{loc}}(V,E)$ for each $\gamma\in\mathcal{F}_{\text{loc}}(U,E)$ and the restriction map
@@
Let $E$ be a finite-dimensional vector space, $U$ and $V$ be open subsets of $[0,\infty)^{m}$ and $\Theta:U\to V$ be a smooth diffeomorphism. Then $\gamma\circ\Theta\in\mathcal{F}_{\text{loc}}(U,E)$ for each $\gamma\in\mathcal{F}_{\text{loc}}(V,E)$ and the map
@@
Let $E$ be a finite-dimensional vector space, $U_{1},...,U_{n}$ be open subsets of $[0,\infty)^{m}$ and $\gamma_{j}\in\mathcal{F}_{\text{loc}}(U_{j},E)$ for $j\in\{1,...,n\}$ such that
@@ 3 Space of $\mathcal{F}$-sections
Let $m,n\in\mathbb{N}$. We assume that $\mathcal{U}$ is a good collection of open subsets of $[0,\infty)^{m}$ and $\left(\mathcal{F}(U,\mathbb{R})\right)_{U\in\mathcal{U}}$ is a family of locally convex spaces suitable for global analysis.
@@
$\Gamma_{\mathcal{F}}(\gamma):=\{\sigma\in\mathcal{F}(M,TN):\pi_{TN}\circ\sigma=\gamma\}$
@@
Hence $\sigma$ make sense. Moreover, we see that $\pi_{TN}\circ\sigma=\gamma$.
@@
Let $\mathcal{U}$ be a good collection of open subsets of $[0,\infty)^{m}$ and $(\mathcal{F}(U,\mathbb{R}))_{U\in\mathcal{U}}$ be a family of locally convex spaces suitable for global analysis. Let $M$ be an $m$-dimensional compact smooth manifold with corners, $N_{1}$ and $N_{2}$ be finite-dimensional smooth manifolds and $\gamma\in\mathcal{F}(M,N_{1})$. If $f:N_{1}\to N_{2}$ is a smooth function, then $Tf\circ\sigma\in\Gamma_{\mathcal{F}}(f\circ\gamma)$ for each $\sigma\in\Gamma_{\mathcal{F}}(\gamma)$. Moreover, the map
@@
$\Gamma_{C}(\gamma)=\{\sigma\in BC(M,TN):\pi_{TN}\circ\sigma=\gamma\}$
@@
is open in $\Gamma_{\mathcal{F}}(\gamma)$ for each open set $V\subseteq TN$
@@
Let $\mathcal{U}$ be a good collection of open subsets of $[0,\infty)^{m}$ and $(\mathcal{F}(U,\mathbb{R}))_{U\in\mathcal{U}}$ be a family of locally convex spaces suitable for global analysis. Let $M$ be a $m$-dimensional compact smooth manifold with corners, $N_{1}$, $N_{2}$ be smooth manifolds $n_{1}$-dimensional and $n_{2}$-dimensional respectively and $\pi_{j}:N_{1}\times N_{2}\to N_{j}$ be the $j$-projection for $j\in\{1,2\}$. If $\gamma_{1}\in\mathcal{F}(M,N_{1})$ and $\gamma_{2}\in\mathcal{F}(M,N_{2})$ then the map
@@
Let $\mathcal{U}$ be a good collection of open subsets of $[0,\infty)^{m}$ and $(\mathcal{F}(U,\mathbb{R}))_{U\in\mathcal{U}}$ be a family of locally convex spaces suitable for global analysis. Let $M_{1}$ and $M_{2}$ be compact smooth manifolds with corners and $N$ be a smooth manifold. If $\Theta:M_{1}\to M_{2}$ is a smooth diffeomorphism, then $\gamma\circ\Theta\in\mathcal{F}(M_{1},N)$ for each $\gamma\in\mathcal{F}(M_{2},N)$. Moreover, the map
@@
Let $\mathcal{U}$ be a good collection of open subsets of $[0,\infty)^{m}$ and $(\mathcal{F}(U,\mathbb{R}))_{U\in\mathcal{U}}$ be a family of locally convex spaces suitable for global analysis. Let $M$ be a compact manifold with corners and $N$ be a smooth manifold. Then the evaluation map
@@ 4 Manifolds of $\mathcal{F}$-maps on compact manifolds
Let $N$ be a smooth manifold and $\pi_{TN}:TN\to N$ its tangent bundle. A local addition is a smooth map ${\Sigma:\Omega\to N}$ defined on a open neightborhood $\Omega\subseteq TN$ of the zero-section ${0_{N}:=\{0_{p}\in T_{p}N:p\in N\}}$ such that
@@
    The image $\Omega^{\prime}:=\big{(}\pi_{TN},\Sigma\big{)}(\Omega)$ is open in $N\times N$ and the map
    $\theta_{N}:\Omega\to\Omega^{\prime},\quad v\mapsto\big{(}\pi_{TN}(v),\Sigma(v)\big{)}$ (4.1)
@@
If a smooth manifold $N$ admits a local addition, then also its tangent manifold $TN$ admits a local addition [3, Lemma A.11]. Moreover, each manifold which admits a local addition also admits a normalized local addition [3, Lemma A.14]. From now we will assume that each local addition is normalized.
@@
Let $\mathcal{U}$ be a good collection of open subsets of $[0,\infty)^{m}$ and $(\mathcal{F}(U,\mathbb{R}))_{U\in\mathcal{U}}$ is a family of locally convex spaces suitable for global analysis. Let $M$ be $m$-dimensional compact manifold with corners and $N$ be a smooth manifold which admits a local addition $\Sigma:\Omega\to N$. Let $\gamma\in\mathcal{F}(M,N)$. We define the set
@@
Let $\mathcal{U}$ be a good collection of open subsets of $[0,\infty)^{m}$ and $(\mathcal{F}(U,\mathbb{R}))_{U\in\mathcal{U}}$ be a family of locally convex spaces suitable for global analysis, then for each compact manifold $M$ with corners and smooth manifold $N$ without boundary which admits a local addition, the set $\mathcal{F}(M,N)$ admits a smooth manifold structure such that the sets ${\mathcal{U}}_{\gamma}$ are open in $\mathcal{F}(M,N)$ for all $\gamma\in\mathcal{F}(M,N)$ and $\Psi_{\gamma}\colon{\mathcal{V}}_{\gamma}\to{\mathcal{U}}_{\gamma}$ is a $C^{\infty}$-diffeomorphism.
@@
Let $\mathcal{U}$ be a good collection of open subsets of $[0,\infty)^{m}$ and $(\mathcal{F}(U,\mathbb{R}))_{U\in\mathcal{U}}$ be a family of locally convex spaces suitable for global analysis. For each compact manifold $M$ with corners and $\mathbb{K}$-analytic manifold $N$ without boundary which admits a $\mathbb{K}$-analytic local addition, the set $\mathcal{F}(M,N)$ admits a $\mathbb{K}$-analytic manifold structure such that the sets ${\mathcal{U}}_{\gamma}$ are open in $\mathcal{F}(M,N)$ for all $\gamma\in\mathcal{F}(M,N)$ and $\Psi_{\gamma}\colon{\mathcal{V}}_{\gamma}\to{\mathcal{U}}_{\gamma}$ is a $C^{\infty}$-diffeomorphism.
@@
Let $\mathcal{U}$ be a good collection of open subsets of $[0,\infty)^{m}$ and $(\mathcal{F}(U,\mathbb{R}))_{U\in\mathcal{U}}$ be a family of locally convex spaces suitable for global analysis. Let $M$ be an $m$-dimensional compact smooth manifold with corners, $N_{1}$ and $N_{2}$ be finite dimensional smooth manifold which admits local addition. If $f:N_{1}\to N_{2}$ is a $smooth$ map, then the map
@@
$F(\sigma):=\Psi_{f\circ\gamma}^{-1}\circ\mathcal{F}(M,f)\circ\Psi_{\gamma}(\sigma)=(\pi_{TN},\Sigma_{2})^{-1}\circ\Big{(}f\circ\gamma,f\circ\Sigma_{1}\circ\sigma\Big{)}$
@@
$d\phi_{2,i}\circ F(\sigma)\circ\varphi_{M,i}^{-1}|_{W_{M,i}^{\prime}}$ $=d\phi_{2,i}\circ(\pi_{TN},\Sigma_{2})^{-1}\circ\Big{(}f\circ\gamma,f\circ\Sigma_{1}\circ\sigma\Big{)}\circ\varphi_{M,i}^{-1}|_{W_{M,i}^{\prime}}$
$=d\phi_{2,i}\circ(\pi_{TN},\Sigma_{2})^{-1}\circ\Big{(}f\circ\gamma\circ\varphi_{M,i}^{-1}|_{W_{M,i}^{\prime}},f\circ\Sigma_{1}\circ\sigma\circ\varphi_{M,i}^{-1}|_{W_{M,i}^{\prime}}\Big{)}.$
@@
$H_{i}(x,y,z):=d\phi_{2,i}\circ(\pi_{TN},\Sigma_{2})^{-1}\circ(\phi_{2,i}^{-1}\left(x\right),f\circ\Sigma_{1}\circ T\phi_{1,i}^{-1}\left(y,z\right)).$
@@
Let $\mathcal{U}$ be a good collection of open subsets of $[0,\infty)^{m}$ and $(\mathcal{F}(U,\mathbb{R}))_{U\in\mathcal{U}}$ be a family of locally convex spaces suitable for global analysis. Let $\mathbb{K}\in\{\mathbb{R},\mathbb{C}\}$, $M$ be an $m$-dimensional compact smooth manifold with corners, $N_{1}$ and $N_{2}$ be $n$-dimensional $\mathbb{K}$-analytic manifolds with $\mathbb{K}$-analytic local additions $(\Omega_{1},\Sigma_{1})$ and $(\Omega_{2},\Sigma_{2})$ respectively. If $f:N_{1}\to N_{2}$ is a $\mathbb{K}$-analytic map, then the map
@@
Let $\mathcal{U}$ be a good collection of open subsets of $[0,\infty)^{m}$ and $(\mathcal{F}(U,\mathbb{R}))_{U\in\mathcal{U}}$ be a family of locally convex spaces suitable for global analysis. Let $M$ be a $m$-dimensional compact smooth manifold with corners, $N_{1}$ and $N_{2}$ be smooth manifolds which admit local additions, and let $\text{pr}_{i}:N_{1}\times N_{2}\to N_{i}$ be the i-th projection where $i\in\{1,2\}$, then the map

### v2

@@ 1 Introduction
For fixed $m,n\in\mathbb{N}$, we consider a $m$-dimensional compact smooth manifold with corners $M$ and $N$ be a $n$-dimensional smooth manifold without boundary. We consider a basis of the topology $\mathcal{U}$ of the set $0,\infty)^{m}$ satisfying suitable properties (see Definition 2.1). Suppose that for each open set $U\in\mathcal{U}$, an integral complete locally convex space $\mathcal{F}(U,\mathbb{R})$ of bounded, continuous real-valued functions are given. Then for each finite-dimensional real vector space $E$, the set of maps $\mathcal{F}(U,E)$ can be defined in a natural way. If certain axioms are satisfied (see Definition 2.5), we say that the family $(\mathcal{F}(U,E))_{U\in\mathcal{U}}$ is suitable for global analysis. As direct consequence of the case where $M$ is a smooth manifold without boundary (see [15]), it can be shown that one can define a locally convex space $\mathcal{F}(M,E)$. Moreover, we can also define the set of mappings $\mathcal{F}(M,N)$ of $N$-valued functions on the manifold with corners $M$.
@@
Let $N$ be a smooth manifold and $\pi_{N}:TN\to N$ its tangent bundle. A local addition is a smooth map $\Sigma:\Omega\to N$} defined on a open neightborhood $\Omega\subseteq TN$ of the zero-section $0_{N}:=\{0_{p}\in T_{p}N:p\in N\}$} such that $\Sigma(0_{p})=p$ for all $p\in N$ and the image $\Omega^{\prime}:=\big(\pi_{N},\Sigma\big)(\Omega)$ is open in $N\times N$ and the map $\theta_{N}:=(\pi_{N},\Sigma):\Omega\to\Omega^{\prime}$ is a $C^{\infty}$-diffeomorphism.
@@
$\Gamma_{\mathcal{F}}(\eta):=\{\sigma\in\mathcal{F}(M,TN):\pi_{N}\circ\sigma=\gamma\}$
@@
Let $M$ be a $m$-dimensional compact smooth manifold with corners, $N_{1}$ and $N_{2}$ be $n$-dimensional smooth manifold which admits local addition $(\Omega_{1},\Sigma_{1})$ and $(\Omega_{2},\Sigma_{2})$ respectively. If $f:N_{1}\to N_{2}$ is a smoothsmooth map, then the map
@@
Let $M$ be a $m$-dimensional compact smooth manifold with corners, $N$ be a $n$-dimensional smooth manifold which admits a local addition and $\pi_{N}:TN\to N$ its tangent bundle. Then the map
$\mathcal{F}(M,\pi_{N}):\mathcal{F}(M,TN)\to\mathcal{F}(M,N),\quad\tau\mapsto\pi_{N}\circ\tau$
@@ 2 Preliminaries
Let $m\in\mathbb{N}$ fixed, a set $\mathcal{U}$ of open subsets of product set $0,\infty)^{m}$ will be called a good collection of open subsets if the following condition are satisfied:
@@
    $\mathcal{U}$ is a basis for the topology of $0,\infty)^{m}$.
@@
    If $U\in\mathcal{U}$ and $K\subseteq U$ is a compact non-empty subset, then there exists $V\in\mathcal{U}$ with compact closure $\overline{V}$ in $0,\infty)^{m}$ such that $K\subseteq V$ and $\overline{V}\subseteq U$.
@@
    If $\phi:U\to V$ is a $C^{\infty}$-diffeomorphism between open subsets $U$ and $V$ of $0,\infty)^{m}$ and $W\in\mathcal{U}$ is a relatively compact subset of $U$, then $\phi(W)\in\mathcal{U}$.
@@
Let $U$ be a open subset of $0,\infty)^{m}$, we write $BC(U,\mathbb{R})$ for the vector space of all bounded continuous functions $f:U\to\mathbb{R}$ endowed with the supremum norm $\lVert\cdot\lVert_{\infty}$.
@@
Let $\mathcal{U}$ be a good collection of open subsets of $0,\infty)^{m}$. For $U\in\mathcal{U}$, the vector subspace $\mathcal{F}(U,\mathbb{R})$ of $BC(U,\mathbb{R})$ will denote a integral complete locally convex space such that the inclusion map $\mathcal{F}(U,\mathbb{R})\to BC(U,\mathbb{R})$ is continuous.
@@
    $f_{*}:\mathcal{F}(U,E)\to\mathcal{F}(V,F),\quad\gamma\mapsto f\circ(\text{id}_{V},\gamma|_{V})$
@@
    Pullback Axiom : Let $U$ be an open subset of $0,\infty)^{m}$ and $V,W\in\mathcal{U}$ such that $W$ has compact closure contained in $U$. Let $\Theta:U\to V$ be a smooth diffeomorphism. Then $\gamma\circ\Theta|_{W}\in\mathcal{F}(W,E)$ for all $\gamma\in\mathcal{F}(V,E)$ and
@@
Following [15, Remark 3.5], if $\mathcal{U}$ is a good collection of open subsets of $0,\infty)^{m}$ and $\left(\mathcal{F}(U,\mathbb{R})\right)_{U\in\mathcal{U}}$ is a family of locally convex space suitable for global analysis, then we have the following results.
@@
which extends functions by 0 is a linear isometry. ∎
@@
$I(t)(x)$ $=$ $\varepsilon_{x}(I(t))=\int_{0}^{1}\varepsilon_{x}(df\circ(\gamma+st\eta,\eta)|_{W})\,ds$
$=$ $\int_{0}^{1}df(\gamma(x)+st\eta(x),\eta(x))\,ds=\frac{f(\gamma(x)+t\eta(x))-f(\gamma(x))}{t}$
@@
Let $U$ be a open subset of $0,\infty)^{m}$ and $E$ be a finite-dimensional real vector space. We let $\mathcal{F}_{\text{loc}}(U,E)$ be the set of all function $\gamma:U\to E$ such that for each $V\in\mathcal{U}$ which is relatively compact in $U$ we have $\gamma|_{V}\in\mathcal{F}(V,E)$.
@@
Let $E$ be a finite-dimensional vector space. If $U$ and $V$ are open subsets of $0,\infty)^{m}$ such that $V\subseteq U$, then $\gamma|_{V}\in\mathcal{F}_{\text{loc}}(V,E)$ for each $\gamma\in\mathcal{F}_{\text{loc}}(U,E)$ and the restriction map
@@
Let $E$ be a finite-dimensional vector space, $U$ and $V$ be open subsets of $0,\infty)^{m}$ and $\Theta:U\to V$ be a smooth diffeomorphism. Then $\gamma\circ\Theta\in\mathcal{F}_{\text{loc}}(U,E)$ for each $\gamma\in\mathcal{F}_{\text{loc}}(V,E)$ and the map
@@
Let $E$ be a finite-dimensional vector space, $U_{1},...,U_{n}$ be open subsets of $0,\infty)^{m}$ and $\gamma_{j}\in\mathcal{F}_{\text{loc}}(U_{j},E)$ for $j\in\{1,...,n\}$ such that
@@ 3 Space of $\mathcal{F}$-sections
Let $m,n\in\mathbb{N}$. We assume that $\mathcal{U}$ is a good collection of open subsets of $0,\infty)^{m}$ and $\left(\mathcal{F}(U,\mathbb{R})\right)_{U\in\mathcal{U}}$ is a family of locally convex spaces suitable for global analysis.
@@
$\Gamma_{\mathcal{F}}(\gamma):=\{\sigma\in\mathcal{F}(M,TN):\pi_{N}\circ\sigma=\gamma\}$
@@
Hence $\sigma$ make sense. Moreover, we see that $\pi_{N}\circ\sigma=\gamma$.
@@
Let $\mathcal{U}$ be a good collection of open subsets of $0,\infty)^{m}$ and $(\mathcal{F}(U,\mathbb{R}))_{U\in\mathcal{U}}$ be a family of locally convex spaces suitable for global analysis. Let $M$ be an $m$-dimensional compact smooth manifold with corners, $N_{1}$ and $N_{2}$ be finite-dimensional smooth manifolds and $\gamma\in\mathcal{F}(M,N_{1})$. If $f:N_{1}\to N_{2}$ is a smooth function, then $Tf\circ\sigma\in\Gamma_{\mathcal{F}}(f\circ\gamma)$ for each $\sigma\in\Gamma_{\mathcal{F}}(\gamma)$. Moreover, the map
@@
$\Gamma_{C}(\gamma)=\{\sigma\in BC(M,TN):\pi_{N}\circ\sigma=\gamma\}$
@@
is open in $\Gamma_{\mathcal{F}}(\gamma)$ for each open set $V\subseteq TN$ .
@@
Let $\mathcal{U}$ be a good collection of open subsets of $0,\infty)^{m}$ and $(\mathcal{F}(U,\mathbb{R}))_{U\in\mathcal{U}}$ be a family of locally convex spaces suitable for global analysis. Let $M$ be a $m$-dimensional compact smooth manifold with corners, $N_{1}$, $N_{2}$ be smooth manifolds $n_{1}$-dimensional and $n_{2}$-dimensional respectively and $\pi_{j}:N_{1}\times N_{2}\to N_{j}$ be the $j$-projection for $j\in\{1,2\}$. If $\gamma_{1}\in\mathcal{F}(M,N_{1})$ and $\gamma_{2}\in\mathcal{F}(M,N_{2})$ then the map
@@
Let $\mathcal{U}$ be a good collection of open subsets of $0,\infty)^{m}$ and $(\mathcal{F}(U,\mathbb{R}))_{U\in\mathcal{U}}$ be a family of locally convex spaces suitable for global analysis. Let $M_{1}$ and $M_{2}$ be compact smooth manifolds with corners and $N$ be a smooth manifold. If $\Theta:M_{1}\to M_{2}$ is a smooth diffeomorphism, then $\gamma\circ\Theta\in\mathcal{F}(M_{1},N)$ for each $\gamma\in\mathcal{F}(M_{2},N)$. Moreover, the map
@@
Let $\mathcal{U}$ be a good collection of open subsets of $0,\infty)^{m}$ and $(\mathcal{F}(U,\mathbb{R}))_{U\in\mathcal{U}}$ be a family of locally convex spaces suitable for global analysis. Let $M$ be a compact manifold with corners and $N$ be a smooth manifold. Then the evaluation map
@@ 4 Manifolds of $\mathcal{F}$-maps on compact manifolds
Let $N$ be a smooth manifold and $\pi_{N}:TN\to N$ its tangent bundle. A local addition is a smooth map $\Sigma:\Omega\to N$} defined on a open neightborhood $\Omega\subseteq TN$ of the zero-section $0_{N}:=\{0_{p}\in T_{p}N:p\in N\}$} such that
@@
    The image $\Omega^{\prime}:=\big(\pi_{N},\Sigma\big)(\Omega)$ is open in $N\times N$ and the map
    $\theta_{N}:\Omega\to\Omega^{\prime},\quad v\mapsto\big(\pi_{N}(v),\Sigma(v)\big)$ (4.1)
@@
Let $N$ be a smooth manifold which admits a local addition. If $\pi_{TN}:T(TN)\to TN$ denotes the tangent bundle of $TN$ and $\kappa:T(TN)\to T(TN)$ its canonical flip, then $T\Sigma\circ\kappa:\tau(T\Omega)\to TN$ it is a local addition on $TN$ [3, Lemma A.11]. Moreover, each manifold which admits a local addition also admits a normalized local addition [3, Lemma A.14]. From now we will assume that each local addition is normalized.
@@
Let $\mathcal{U}$ be a good collection of open subsets of $0,\infty)^{m}$ and $(\mathcal{F}(U,\mathbb{R}))_{U\in\mathcal{U}}$ is a family of locally convex spaces suitable for global analysis. Let $M$ be $m$-dimensional compact manifold with corners and $N$ be a smooth manifold which admits a local addition $\Sigma:\Omega\to N$. Let $\gamma\in\mathcal{F}(M,N)$. We define the set
@@
Let $\mathcal{U}$ be a good collection of open subsets of $0,\infty)^{m}$ and $(\mathcal{F}(U,\mathbb{R}))_{U\in\mathcal{U}}$ be a family of locally convex spaces suitable for global analysis, then for each compact manifold $M$ with corners and smooth manifold $N$ without boundary which admits a local addition, the set $\mathcal{F}(M,N)$ admits a smooth manifold structure such that the sets ${\mathcal{U}}_{\gamma}$ are open in $\mathcal{F}(M,N)$ for all $\gamma\in\mathcal{F}(M,N)$ and $\Psi_{\gamma}\colon{\mathcal{V}}_{\gamma}\to{\mathcal{U}}_{\gamma}$ is a $C^{\infty}$-diffeomorphism.
@@
Let $\mathcal{U}$ be a good collection of open subsets of $0,\infty)^{m}$ and $(\mathcal{F}(U,\mathbb{R}))_{U\in\mathcal{U}}$ be a family of locally convex spaces suitable for global analysis. For each compact manifold $M$ with corners and $\mathbb{K}$-analytic manifold $N$ without boundary which admits a $\mathbb{K}$-analytic local addition, the set $\mathcal{F}(M,N)$ admits a $\mathbb{K}$-analytic manifold structure such that the sets ${\mathcal{U}}_{\gamma}$ are open in $\mathcal{F}(M,N)$ for all $\gamma\in\mathcal{F}(M,N)$ and $\Psi_{\gamma}\colon{\mathcal{V}}_{\gamma}\to{\mathcal{U}}_{\gamma}$ is a $C^{\infty}$-diffeomorphism.
@@
Let $\mathcal{U}$ be a good collection of open subsets of $0,\infty)^{m}$ and $(\mathcal{F}(U,\mathbb{R}))_{U\in\mathcal{U}}$ be a family of locally convex spaces suitable for global analysis. Let $M$ be an $m$-dimensional compact smooth manifold with corners, $N_{1}$ and $N_{2}$ be finite dimensional smooth manifold which admits local addition. If $f:N_{1}\to N_{2}$ is a smoothsmooth map, then the map
@@
$F(\sigma):=\Psi_{f\circ\gamma}^{-1}\circ\mathcal{F}(M,f)\circ\Psi_{\gamma}(\sigma)=(\pi_{N},\Sigma_{2})^{-1}\circ\Big(f\circ\gamma,f\circ\Sigma_{1}\circ\sigma\Big)$
@@
$d\phi_{2,i}\circ F(\sigma)\circ\varphi_{M,i}^{-1}|_{W_{M,i}^{\prime}}$ $=d\phi_{2,i}\circ(\pi_{N},\Sigma_{2})^{-1}\circ\Big(f\circ\gamma,f\circ\Sigma_{1}\circ\sigma\Big)\circ\varphi_{M,i}^{-1}|_{W_{M,i}^{\prime}}$
$=d\phi_{2,i}\circ(\pi_{N},\Sigma_{2})^{-1}\circ\Big(f\circ\gamma\circ\varphi_{M,i}^{-1}|_{W_{M,i}^{\prime}},f\circ\Sigma_{1}\circ\sigma\circ\varphi_{M,i}^{-1}|_{W_{M,i}^{\prime}}\Big).$
@@
$H_{i}(x,y,z):=d\phi_{2,i}\circ(\pi_{N},\Sigma_{2})^{-1}\circ(\phi_{2,i}^{-1}\left(x\right),f\circ\Sigma_{1}\circ T\phi_{1,i}^{-1}\left(y,z\right)).$
@@
Let $\mathcal{U}$ be a good collection of open subsets of $0,\infty)^{m}$ and $(\mathcal{F}(U,\mathbb{R}))_{U\in\mathcal{U}}$ be a family of locally convex spaces suitable for global analysis. Let $\mathbb{K}\in\{\mathbb{R},\mathbb{C}\}$, $M$ be an $m$-dimensional compact smooth manifold with corners, $N_{1}$ and $N_{2}$ be $n$-dimensional $\mathbb{K}$-analytic manifolds with $\mathbb{K}$-analytic local additions $(\Omega_{1},\Sigma_{1})$ and $(\Omega_{2},\Sigma_{2})$ respectively. If $f:N_{1}\to N_{2}$ is a $\mathbb{K}$-analytic map, then the map
@@
Let $\mathcal{U}$ be a good collection of open subsets of $0,\infty)^{m}$ and $(\mathcal{F}(U,\mathbb{R}))_{U\in\mathcal{U}}$ be a family of locally convex spaces suitable for global analysis. Let $M$ be a $m$-dimensional compact smooth manifold with corners, $N_{1}$ and $N_{2}$ be smooth manifolds which admit local additions, and let $\text{pr}_{i}:N_{1}\times N_{2}\to N_{i}$ be the i-th projection where $i\in\{1,2\}$, then the map